1.(sint)/t


2.(1 - cos t)/t


3.(e ^ t - 1)/t

1
Expert's answer
2022-03-31T17:07:48-0400

Suppose that the task is to compute the limits of these functions as "t\rightarrow0".

  1. "\lim_{t\rightarrow 0}\frac{sin\, t}{t}." It is a well-known limit. It is possible to compute this limit by using geometric considerations or by using L'Hôpital's rule. Namely, "(sin\,t)'=cos\,t" and "t'=1", where "'" denotes the derivative. Therefore, "\lim_{t\rightarrow 0}\frac{sin\, t}{t}=\lim_{t\rightarrow 0}\frac{cos\, t}{1}=1"
  2. "\lim_{t\rightarrow 0}\frac{1-cos\, t}{t}". Using trigonometric formulae, one receives: "1-cos\,t=2\,sin^2\,\frac{t}{2}". We use the limit from the previous exercise and get: "\lim_{t\rightarrow 0}\frac{1-cos\, t}{t}=\lim_{t\rightarrow 0}2\frac{sin^2\, \frac{t}{2}}{\frac{t^2}{4}}\frac{t}4=0."
  3. "\lim_{t\rightarrow 0}\frac{e^t-1}{t}." It is also a well-known limit. It is possible to compute the limit using different considerations. We use L'Hôpital's rule: "(e^t-1)'=e^t,\quad t'=1". Therefore: "\lim_{t\rightarrow 0}\frac{e^t-1}{t}.=\lim_{t\rightarrow 0}e^t=1."